What is the equivalent resistance between points a


please show work
Answer #1

10 and 10 ohms in parallel:

Rp = R1 * R2 / ( R1 + R2 )

     = 10 * 10 / ( 10 + 10 )

    = 5 ohms

5 ohms and 5 ohms in series:

Rs = R1 + R2

    = 5 + 5

   = 10 ohms

10 and 10 ohms in parallel:

Rp = R1 * R2 / ( R1 + R2 )

     = 10 * 10 / ( 10 + 10 )

    = 5 ohms

5 ohms and 20 ohms in series:

Rab = R1 + R2

    = 5 + 20

= 25 ohms